CF/(1+r) - INV vestment the questions 1. Based on the following information, answer the questions, 1) Facebook issued 10-yearplease help with part 2, # 1) and 2)
Answer #1

1.
Price using constant growth dividend model
=D1/(r-g)
=8/(20%-0%)
=40

2.
=6/1.2+7/1.2^2+40/1.2^2
=37.63888889